# Low-Loss Space in Neural Networks is Continuous and Fully Connected

### v1

@@ 1 Introduction
The training of neural networks is often understood as the process of optimizing a non-convex loss function to locate local or global minima [1]. Previous studies have visualized the loss function, referred to as the loss landscape, suggesting that minima are points, as demonstrated in Figure 1 of Foret et al. [2], Figure 1 and Figure 4 of Li et al. [3]. By using different random seeds, it is possible to find different minima, which suggests that the total low-loss space may be a collection of isolated points.
However, these loss landscape visualizations are limited to low-dimensional projections, far smaller than the full parameter space, typically represented as $\mathbb{R}^{D}$, where $D$ denotes the number of trainable parameters. Recent theoretical advancements indicate that low-loss points should form connected regions rather than isolated points [4, 5, 6], implying the existence of a path where the loss remains consistently low. Empirical studies support this, demonstrating the feasibility of constructing paths between distinct low-loss points by slightly perturbing the linear interpolation between two low-loss points [7, 8].
In this study, we investigate the geometry of the low-loss space, defined as the set of all model parameters in $\mathbb{R}^{D}$ that yield a loss below a specified threshold. Unlike previous studies which only constructs a low-loss path between two low-loss points [7, 8], our algorithms are able to find low-loss paths in the full $\mathbb{R}^{D}$ space. We demonstrate that the low-loss space is not only continuous but also fully-connected as a single subspace within $\mathbb{R}^{D}$. By continuous, we mean that directions exist along which a low-loss model can be moved without leaving the low-loss space [9]. By fully-connected, we imply the entire low-loss space can be traversed without exiting this space.
The above statements are supported by our experimental results. We design two algorithms: (1) one algorithm to construct low-loss paths between two models and (2) one algorithm to construct low-loss paths from a low-loss point toward the origin. We test our algorithms on LeNet5 [10], ResNet18 [11], and Compact Convolutional Transformer (CCT7) [12]. Results indicate that such paths consistently exist regardless of the low-loss points chosen, providing evidence that the low-loss space is both continuous and connected. Moreover, since these points represent models derived from independent training processes with different random seeds yet remain interconnected, we conclude that the low-loss space constitutes a fully-connected space.
@@
The second observation involves visualization through principal component analysis (PCA). We apply PCA to project the low-loss space into two dimensions, offering a novel perspective on the representation of features captured by individual layers. The third observation focuses on the relationship between model generalization and weight variance—typically regulated through regularization techniques in existing studies [13, 14]. Our results suggest that it is possible to further improve the generalization of the state-of-the-art ResNet18 model, on top of existing data augmentation techniques [15] and training hyperparameters fine-tuning.
@@ 2 Related Work
Existing theoretical studies have shown that the low-loss region of neural network is continuous. Watanabe’s Singular Learning Theory (SLT) [16, 17, 18, 19], which is grounded in Bayesian learning, suggests that neural networks are singular models. The loss landscape of a singular model is a valley, with minima creating a continuous path at the bottom of the valley, as illustrated in Figure 3.1 of Carroll [5] and discussed in informal articles [20]. Additional studies reach similar conclusions by showing the existence of nearby low-loss points adjacent to a given minimum [4, 21].
Empirically, several algorithms have been proposed to construct continuous paths between two low-loss points. Among them, one of the most prominent is the Automated Nudged Elastic Band (AutoNEB) method [8]. AutoNEB begins by constructing a linear path between two models and then iteratively bends the path to reduce the loss at intermediate points. However, this method has several limitations. Most notably, it does not guarantee the discovery of a continuous low-loss path, as evidenced in Figure 3 of [8] [^1] In contrast, our algorithm demonstrates significantly more consistent low-loss paths.. This limitation arises because the midpoint—typically the unweighted average of two models—often exhibits poor accuracy [22], thereby acting as a barrier along the path. In favorable scenarios, overcoming this barrier requires significant additional computational cost (as shown in Figure 5 of [8]); in less favorable cases, AutoNEB fails to find a viable path.
Other empirical efforts include studies that explore polynomial interpolation between low-loss points to approximate low-loss paths [7], as well as those that investigate the barriers on the interpolation [23, 24].
@@ 3.1 Terminology and Definitions
Unlike the concept of local minima on the loss surface, where the location of a minimum is found by a stochastic process, the low-loss space is a well-defined region for a given model architecture $\mathcal{M}$ and dataset $\mathcal{D}$, denoted as $\text{LLS}(\mathcal{M},\mathcal{D})=S_{L\leq L_{\text{thres}}}$. In principle, this space could be identified by evaluating the loss function at all points in $\mathbb{R}^{D}$; however, this is infeasible in practice, as modern neural network architectures can have billions of parameters, making exhaustive evaluation computationally prohibitive.
@@
In typical model training procedures, the parameters of linear layers, convolutional layers, and transformer layers are initialized as a distribution with a specific variance and mean. The mean is usually set to zero, while the variance is determined by one of the following methods: LeCun [25], Xavier [26], or Kaiming [27]. The number of trainable parameters also affects initialization, as all above methods compute variance based on the number of trainable parameters in the given layer. This relationship is mathematically expressed in Equation 3.
@@
Our experimental results indicate that models trained with identical hyperparameters tend to be positioned on variance spheres that are close to each other. However, this observation is not universally applicable across all model architectures and must be empirically validated before applying our algorithms. Our experimental validation of this relationship for ResNet18 is provided in Appendix A.1. This relationship is mathematically expressed in Equation 3.1.
$\theta_{n}$ $=\text{Train}^{n}(\theta_{0},\mathbf{x},\mathcal{D}),\quad\theta_{n}^{\prime}=\text{Train}^{n}(\theta_{0}^{\prime},\mathbf{x},\mathcal{D})$
$\implies\operatorname{Var}(\theta_{n})\approx\operatorname{Var}(\theta_{n}^{\prime})$ (4)
@@
This equation holds due to the following reasons: (1) during initialization, layer weights (excluding normalization layers) are set to a distribution with zero mean; (2) after training iterations, the mean of layer weights deviates from zero. This deviation is small and can be ignored since it does not influence our algorithms.
@@ 3.2 LLPF Algorithm
Algorithm 1 LLPF_M2M: Algorithm for constructing a low-loss path between two models. The major steps of this algorithm is also illustrated in the geometry plot in Figure 1, which uses consistent point notations as this algorithm.
@@
       $M_{3}$ = Trainr($M_{2}$, $\mathbf{x}$, $\mathcal{B}$) $\triangleright$ Corresponding to $\overrightarrow{M_{2}M_{3}}$. Training $n$ steps with hyper-parameters $\mathbf{x}$
@@
We propose the Low-Loss Path Finding (LLPF) algorithm to explore the subspace $\mathcal{S}_{L<L_{\mathrm{thres}}}\subset\mathbb{R}^{D}$, which comprises all model parameters yielding a loss below a specified threshold. LLPF is composed of two complementary components. The first, referred to as the model-to-model algorithm (LLPF_M2M), constructs a path between two low-loss models located on variance spheres with a small variance difference—that is, satisfying the condition in Equation 3.1. This procedure is detailed in Algorithm 1. The second component, the model-to-origin algorithm (LLPF_M2O), generates a path from a low-loss model toward the origin of $\mathbb{R}^{D}$, as described in Algorithm 2. These components serve distinct purposes: Algorithm 1 enables exploration within the same variance sphere, while Algorithm 2 enables exploration across variance spheres.
The operation of Algorithm 1 is illustrated geometrically in Figure 1, using the notation of the algorithm, including points $P_{0}$, $P_{1}$, $M_{1}$, $M_{2}$, and $M_{3}$. The algorithm initiates by moving the starting point $P_{0}$ slightly toward the destination $D$ to produce an intermediate model $M_{1}$ (Line 5). This Move operation performs a weighted average of model parameters, with weights determined by hyperparameters step_f and step_a. Next, the VarianceCorrection step projects $M_{1}$ back as $M_{2}$ onto the variance sphere of $P_{0}$ to counteract the vanishing variance problem, a phenomenon wherein averaging uncorrelated neural networks leads to reduced parameter variance, thereby hindering subsequent training efforts (see [28]). At Line 7, $M_{2}$ undergoes $r$ training iterations to reduce the loss below threshold, yielding $M_{3}$. Finally, $M_{3}$ is projected back onto the variance sphere to produce $P_{1}$, which serves as the new starting point for the next iteration. By repeating this process for $T$ iterations with appropriately chosen hyperparameters, the algorithm generates a sequence $\{P_{i}\}$ that forms a continuous low-loss path from $P_{0}$ to $P_{T}$. The final point $P_{T}$ should reach the target model $D$, thereby completing the model-to-model path construction.
The mechanism of Algorithm 2 is similar to that of Algorithm 1, with two key differences: the removal of the VarianceCorrection step and the inclusion of an AngleConformal step. The AngleConformal step is introduced to adaptively reduce the learning rate as the variance decreases. This adjustment is necessary because applying the learning rate appropriate for a large variance sphere [^2] A large variance sphere refers to a variance sphere with a high variance value. to a small variance sphere often causes the model to deviate from the low-loss path. Although Algorithm 2 can theoretically be applied in reverse (to move a model toward a larger variance sphere) we observe that this process frequently suffers from exploding gradients problem [29]. Consequently, the reverse direction is omitted.
@@
For Algorithm 1, the optimizer should exclude regularization techniques such as weight decay, as the constraint of remaining within a single variance sphere already ensures that the variance of the layer weights is preserved. Additionally, momentum is generally avoided in this setting. While momentum is usually used to stabilize long-distance training trajectories [^3] From a randomly initialized model to a low-loss model., it is less suitable here, as the update steps in Algorithm 1 are small and localized.
@@
8:       $P_{i+1}$ = Trainr($N$, $\mathbf{x}$, $\mathcal{B}$) $\triangleright$ Training $n$ steps with hyper-parameters $\mathbf{x}$
@@
Algorithm 1 has two prerequisites. First, corresponding layers of the input models must lie on nearby variance spheres, as defined in Equation 3.1; a counterexample is VGG16 (without batch normalization) on CIFAR-10, where both Algorithm 1 and Algorithm 2 are needed to explore the low-loss space. Second, shared with Algorithm 2, input models should lie in flat low-loss regions, similar to the structure shown in Figure 1 (right) of [2]. Empirically, we find that starting from sharp minima often causes the loss to gradually increase during path construction. From our previous experiments, ResNet18 models trained with the Adam optimizer on CIFAR-10 tend to converge to sharp minima, preventing Algorithm 1 from constructing a valid low-loss path [^4] This issue can be addressed using the configuration in Table 3.. While sharpness is hard to measure precisely, we find that minima with near-zero training loss are generally flat and suitable as starting points.
@@ 4 Experimental Evaluation of Low-Loss Path Finding Algorithms
This section presents the results of applying Algorithm 1 and Algorithm 2 to explore the LLS(LeNet5, MNIST), LLS(ResNet18,CIFAR10) and LLS(CCT7,CIFAR10), as shown in Figure 3 and Figure 3. To ensure consistency of the results, we repeat each experiment 200, 40, and 10 times, respectively, using different low-loss models trained with identical hyperparameters. Full hyperparameter settings are listed in Appendix A.4.
In Figure 3, we assess path validity using two criteria: (1) the training loss must remain below $L_{\mathrm{thres}}$ (first row); and (2) the distance between $P_{i}$ and $D$ should gradually converge to zero (second row). Minor deviations are acceptable due to Equation 3.1, which states that the starting point $P_{0}$ and destination $D$ are approximately, but not exactly, on the same variance sphere. In our results, all training losses remain below 0.1, and the final layer-wise distances gradually converges below $10^{-1}$, indicating that the low-loss paths in Figure 3 are correctly constructed.
The third row of Figure 3 shows test accuracy and loss along the paths. However, we observe that Algorithm 1 does not guarantee low testing loss, as seen in the ResNet18 case, where test loss increases despite low training loss. This discrepancy arises because the training and testing datasets are different, and define different low-loss spaces, as per Equation 1. Algorithm 1 only ensures that intermediate points fall within the low-loss space defined by the training dataset, not the testing dataset.
Figure 2: Results of Algorithm 1. Each column corresponds to a different model and dataset: LeNet5@MNIST (first column), ResNet18@CIFAR10 (second column), and CCT7@CIFAR10 (third column). The first row shows the training loss; the second row shows the layer-wise L2 distance between $P_{i}$ and the destination point $D$; and the third row shows the test loss and test accuracy. Legends for the second row are omitted due to the large number of layers. Each experiment is repeated with different starting and destination models, trained from different random seeds. The curves show the mean across repetitions, and the shaded regions indicate the standard deviation. Both the curves and the shaded boundaries are smoothed using a moving average with a window size of 10.
@@
Figure 3 highlights additional insights. First, the training loss for LeNet5 increases because all layers are moved toward the origin, and all layer variances decrease. In contrast, ResNet18 avoids this issue as its normalization layers are excluded from movement. Second, we observe a peak in test accuracy or a minimum in test loss along the path toward the origin. This phenomenon aligns with the intuition behind model regularization techniques, such as L2 norm and weight decay, which encourage models to stay closer to the origin to improve model generalization [30, 31]. However, traditional regularization relies on discrete hyperparameter tuning. In contrast, Algorithm 2 reveals a smoother relationship between model generalization and distance to the origin. Moreover, it requires less hyperparameter tuning and is applied post-training, offering a more efficient alternative to conducting multiple rounds of training with different regularization settings.
@@
We also compare LLPF with AutoNEB [8], which bends linear interpolation between low-loss models to form a low-loss path. Our proposed LLPF algorithms differ from AutoNEB in several key aspects. First, unlike AutoNEB, which often finds paths with low testing loss, LLPF does not guarantee this, highlighting an inherent difference in the resulting low-loss paths. Second, LLPF-generated paths exhibit consistent properties, such as test loss and L2 distance to the origin, across different simulations, as shown by the narrow shaded regions in Figure 3 and Figure 3, whereas AutoNEB paths are often inconsistent, as shown in Figure 7 of  [8]. Third, AutoNEB does not explore regions near the origin, while Algorithm 2 explicitly targets this space.

### v3

@@ 1 Introduction
The training of neural networks is often understood as the process of optimizing a non-convex loss function to locate local or global minima 10.5555/3305890.3305950 . Previous studies have visualized the loss function, referred to as the loss landscape, suggesting that minima are points, as demonstrated in Figure 1 of Foret et al. Sharpness_aware_Minimization_for_Efficiently_Improving_Generalization , Figure 1 and Figure 4 of Li et al. Visualizing_the_Loss_Landscape_of_Neural_Nets . By using different random seeds, it is possible to find different minima, which suggests that the total low-loss space may be a collection of isolated points.
However, these loss landscape visualizations are limited to low-dimensional projections, far smaller than the full parameter space, typically represented as $\mathbb{R}^{D}$, where $D$ denotes the number of trainable parameters. Recent theoretical advancements indicate that low-loss points should form connected regions rather than isolated points 10.5555/3546258.3546401 ; carroll2021phase ; freeman2017topologygeometryhalfrectifiednetwork , implying the existence of a path where the loss remains consistently low. Empirical studies support this, demonstrating the feasibility of constructing paths between distinct low-loss points by slightly perturbing the linear interpolation between two low-loss points NEURIPS2018_be3087e7 ; pmlr-v80-draxler18a .
In this study, we investigate the geometry of the low-loss space, defined as the set of all model parameters in $\mathbb{R}^{D}$ that yield a loss below a specified threshold. Unlike previous studies which only constructs a low-loss path between two low-loss points NEURIPS2018_be3087e7 ; pmlr-v80-draxler18a , our algorithms are able to find low-loss paths in the full $\mathbb{R}^{D}$ space. We demonstrate that the low-loss space is not only continuous but also fully-connected as a single subspace within $\mathbb{R}^{D}$. By continuous, we mean that directions exist along which a low-loss model can be moved without leaving the low-loss space 10.1007/978-1-4757-6848-0 . By fully-connected, we mean the entire low-loss space can be traversed without exiting this space.
The above statements are supported by our experimental results. We design two algorithms: (1) one algorithm to construct low-loss paths between two models and (2) one algorithm to construct low-loss paths from a low-loss point toward the origin. We test our algorithms on LeNet5 LeNet , ResNet18 he2015deep , and Compact Convolutional Transformer (CCT7) hassani2022escapingbigdataparadigm . Our results indicate that such paths consistently exist regardless of the low-loss points chosen, providing evidence that the low-loss space is both continuous and connected. Moreover, since these points represent models derived from independent training processes with different random seeds yet remain interconnected, we conclude that the low-loss space constitutes a fully-connected space. Compared to prior empirical work pmlr-v80-draxler18a , our algorithms also yield paths with lower maximum loss, offering stronger quantitative support for the connectedness of the low-loss region.
@@
The second observation involves visualization through principal component analysis (PCA). We apply PCA to project the low-loss space into two dimensions, offering a novel perspective on the representation of features captured by individual layers. The third observation focuses on the relationship between model generalization and weight variance—typically regulated through regularization techniques in existing studies zhang2018three ; loshchilov2018decoupled . Our results suggest that it is possible to further improve the generalization of the state-of-the-art ResNet18 model, on top of existing data augmentation techniques MUMUNI2022100258 and training hyperparameters fine-tuning.
@@ 2 Related Work
Existing theoretical studies have shown that the low-loss region of neural network is continuous. Watanabe’s Singular Learning Theory (SLT) 4233934 ; Watanabe_2009 ; 10.5555/2567709.2502609 ; Watanabe2018 , which is grounded in Bayesian learning, suggests that neural networks are singular models. The loss landscape of a singular model is a valley, with minima creating a continuous path at the bottom of the valley, as illustrated in Figure 3.1 of Carroll carroll2021phase and discussed in informal articles lesswrongNeuralNetworks . Additional studies reach similar conclusions by showing the existence of nearby low-loss points adjacent to a given minimum 10.5555/3546258.3546401 ; 155333 .
Empirically, several algorithms have been proposed to construct continuous paths between two low-loss points. Among them, one of the most prominent is the Automated Nudged Elastic Band (AutoNEB) method pmlr-v80-draxler18a . AutoNEB begins by constructing a linear path between two models and then iteratively bends the path to reduce the loss at intermediate points. However, this method has several limitations. Most notably, it does not guarantee the discovery of a continuous low-loss path, as evidenced in Figure 3 of pmlr-v80-draxler18a  [^1] In contrast, our algorithm demonstrates significantly more consistent low-loss paths.. This limitation arises because the midpoint—typically the unweighted average of two models—often exhibits poor accuracy ju2017relativeperformanceensemblemethods , thereby acting as a barrier along the path. In favorable scenarios, overcoming this barrier requires significant additional computational cost (as shown in Figure 5 of pmlr-v80-draxler18a ); in less favorable cases, AutoNEB fails to find a viable path.
Other empirical efforts include studies that explore polynomial interpolation between low-loss points to approximate low-loss paths NEURIPS2018_be3087e7 , as well as those that investigate the barriers on the interpolation vlaar2022linearinterpolationneuralnetwork ; goodfellow2015qualitativelycharacterizingneuralnetwork .
@@ 3.1 Terminology and Definitions
Unlike local minima, which are discovered through stochastic processes, the low-loss space is a well-defined region for a given model architecture $\mathcal{M}$ and dataset $\mathcal{D}$, denoted as $\text{LLS}(\mathcal{M},\mathcal{D})=S_{L\leq L_{\text{thres}}}$. While it is theoretically possible to identify this space by evaluating the loss across $\mathbb{R}^{D}$, such exhaustive search is infeasible due to the high dimensionality of modern neural networks.
@@
In typical model training procedures, the parameters of linear layers, convolutional layers, and transformer layers are initialized as a distribution with a specific variance and mean. The mean is usually set to zero, while the variance is determined by one of the following methods: LeCun 10.5555/645754.668382 , Xavier pmlr-v9-glorot10a , or Kaiming 7410480 . The number of trainable parameters also affects initialization, as all above methods compute variance based on the number of trainable parameters in the given layer. This relationship is mathematically expressed in Equation 3.
@@
Our experimental results indicate that models trained with identical hyperparameters tend to be positioned on variance spheres that are close to each other. However, this observation is not universally applicable across all model architectures and must be empirically validated before applying our algorithms. Our experimental validation of this relationship for ResNet18 is provided in Appendix A.1. This relationship is mathematically expressed in Equation 4.
$\theta_{n}$ $=\text{Train}^{n}(\theta_{0},\mathbf{x},\mathcal{D}),\quad\theta_{n}^{\prime}=\text{Train}^{n}(\theta_{0}^{\prime},\mathbf{x},\mathcal{D})\implies\operatorname{Var}(\theta_{n})\approx\operatorname{Var}(\theta_{n}^{\prime})$ (4)
@@
This equation holds due to the following reasons: (1) during initialization, layer weights (excluding normalization layers) are set to a distribution with zero mean; (2) after training iterations, the mean of layer weights deviates from zero. This deviation is small and can be ignored since it does not influence our algorithms. We justify this approximation empirically in Appendix A.2.
@@ 3.2 LLPF Algorithm
Algorithm 1 LLPF_M2M: Algorithm for constructing a low-loss path between two models. The major steps of this algorithm are also illustrated in the geometry plot in Figure 1, which uses consistent point notations as this algorithm.
@@
       $M_{3}$ = Trainr($M_{2}$, $\mathbf{x}$, $\mathcal{B}$) $\triangleright$ Corresponding to $\overrightarrow{M_{2}M_{3}}$. Training $r$ steps with hyper-parameters $\mathbf{x}$
@@
We propose the Low-Loss Path Finding (LLPF) algorithm to explore the subspace $\mathcal{S}_{L<L_{\mathrm{thres}}}\subset\mathbb{R}^{D}$, which comprises all model parameters yielding a loss below a specified threshold. LLPF is composed of two complementary components. The first, referred to as the model-to-model algorithm (LLPF_M2M), constructs a path between two low-loss models located on variance spheres with a small variance difference—that is, satisfying the condition in Equation 4. This procedure is detailed in Algorithm 1. The second component, the model-to-origin algorithm (LLPF_M2O), generates a path from a low-loss model toward the origin of $\mathbb{R}^{D}$, as described in Algorithm 2. These components serve distinct purposes: Algorithm 1 enables exploration within the same variance sphere, while Algorithm 2 enables exploration across variance spheres.
The geometric intuition behind Algorithm 1 is illustrated in Figure 1, using the notation of the algorithm, including points $P_{0}$, $P_{1}$, $M_{1}$, $M_{2}$, and $M_{3}$. The algorithm starts by moving the starting $P_{0}$ slightly toward the destination $D$ to obtain $M_{1}$ via a weighted average(Move, Line 5), controlled by hyperparameters step_f and step_a. Next, the VarianceCorrection step projects $M_{1}$ back as $M_{2}$ onto the variance sphere of $P_{0}$ to counteract the vanishing variance problem, a phenomenon wherein averaging uncorrelated neural networks leads to reduced parameter variance, thereby hindering subsequent training efforts (see tian2024vanishingvarianceproblemfully ). $M_{2}$ is refined through $r$ training steps to reduce loss below threshold, yielding $M_{3}$, which is again projected back to the sphere to produce $P_{1}$., which serves as the new starting point for the next iteration. Repeating this process for $T$ iterations yields a continuous low-loss path $\{P_{i}\}$ from $P_{0}$ to $P_{T}\approx D$.
The mechanism of Algorithm 2 is similar to that of Algorithm 1, with two key differences: the removal of the VarianceCorrection step and the inclusion of an AngleConformal step. The AngleConformal step is introduced to adaptively reduce the learning rate as the variance decreases. This adjustment is necessary because applying the learning rate appropriate for a large variance sphere [^2] A large variance sphere refers to a variance sphere with a high variance value. to a small variance sphere often causes the model to deviate from the low-loss path. Although Algorithm 2 can theoretically be applied in reverse (to move a model toward a larger variance sphere) we observe that this process frequently suffers from exploding gradients problem 10.5555/3326943.3326997 . Consequently, the reverse direction is omitted.
@@
For both algorithms, two hyperparameters directly influence the quality of the low-loss paths. The first is $step_{a}$ and $step_{f}$, which control the continuity of the path—smaller values lead to smaller distances between consecutive points $P_{i}$ and $P_{i+1}$. The second is $r$, which determines how low the training loss can go—larger values correspond to more training iterations and potentially lower loss. However, higher-quality paths come at the cost of increased computational workload, so the computation budget must be considered.
@@
For Algorithm 1, the optimizer should exclude regularization techniques such as weight decay, as the constraint of remaining within a single variance sphere already ensures that the variance of the layer weights is preserved. Additionally, momentum is generally avoided in this setting. While momentum is usually used to stabilize long-distance training trajectories, it is less suitable here, as the update steps in Algorithm 1 are small and localized.
@@
8:       $P_{i+1}$ = Trainr($N$, $\mathbf{x}$, $\mathcal{B}$) $\triangleright$ Training $r$ steps with hyper-parameters $\mathbf{x}$
@@
Algorithm 1 has two prerequisites. First, corresponding layers of the input models must lie on nearby variance spheres, as defined in Equation 4; a counterexample is VGG16 (without batch normalization) on CIFAR-10, where both Algorithm 1 and Algorithm 2 are needed to explore the low-loss space. Second, shared with Algorithm 2, input models should lie in flat low-loss regions, similar to the structure shown in Figure 1 (right) of Sharpness_aware_Minimization_for_Efficiently_Improving_Generalization . Empirically, we find that starting from sharp minima can cause our algorithm fail to find feasible paths. This is likely because sharp minima are surrounded by many local minima, increasing the chance of our algorithm getting trapped. In our experiments, ResNet18 models trained with the Adam optimizer on CIFAR-10 tend to converge to sharp minima, preventing Algorithm 1 from constructing a valid low-loss path [^3] This issue can be addressed using the configuration in Table 3.. While sharpness is hard to measure precisely, we find that minima with near-zero training loss are generally flat and suitable as starting points.
@@ 4 Experimental Evaluation of Low-Loss Path Finding Algorithms
This section presents the results of applying Algorithm 1 and Algorithm 2 to explore the LLS(LeNet5, MNIST 6296535 ), LLS(ResNet18,CIFAR10 cifar10_dataset ) and LLS(CCT7,CIFAR10), as shown in Figure 3 and Figure 3. To ensure consistency of the results, we repeat each experiment 200, 40, and 10 times, respectively, using different low-loss models trained with identical hyperparameters. Full hyperparameter settings are listed in Appendix A.6.
In Figure 3, we assess path validity using three criteria: (1) the training loss must remain below $L_{\mathrm{thres}}$ (first row); (2) the distance between $P_{i}$ and $D$ should gradually converge to zero (second row), with minor deviations allowed due to Equation 4, which states that the starting point $P_{0}$ and destination $D$ lie approximately—but not exactly—on the same variance sphere; and (3) the final testing loss should converge to the testing loss at iteration 0, since the final point is expected to converge to $D$, which should exhibit similar generalization behavior as $P_{0}$. In our results, all training losses remain below 0.1, and the final layer-wise distances gradually converges below $10^{-1}$.indicating that the low-loss paths in Figure 3 are correctly constructed.
The third row of Figure 3 shows test accuracy and loss along the paths. However, we observe that Algorithm 1 does not guarantee low testing loss, as seen in the ResNet18 case, where test loss increases despite low training loss. This is because the algorithm operates within the training-defined low-loss space (Equation 1), which is not align with the testing dataset.
Figure 2: Results of Algorithm 1. Each column corresponds to a different model and dataset: LeNet5@MNIST (first column), ResNet18@CIFAR10 (second column), and CCT7@CIFAR10 (third column). The first row shows the training loss; the second row shows the layer-wise L2 distance between $P_{i}$ and the destination point $D$; and the third row shows the test loss and test accuracy of the whole testing dataset. Legends for the second row are omitted due to the large number of layers. Each experiment is repeated with different starting and destination models, trained from different random seeds. The curves show the mean across repetitions, and the shaded regions indicate the standard deviation. Both the curves and the shaded boundaries are smoothed using a moving average with a window size of 10.
@@
Figure 3 highlights additional insights. First, LeNet5 shows increasing training loss as all layers move toward the origin and their variances decrease. In contrast, ResNet18 avoids this issue as its normalization layers are excluded from movement. Second, we observe a peak in test accuracy or a minimum in test loss along the path toward the origin. This phenomenon aligns with the intuition behind model regularization techniques, such as L2 norm and weight decay, which encourage models to stay closer to the origin to improve model generalization 10.5555/3495724.3496126 ; NIPS1991_8eefcfdf . Unlike traditional methods requiring hyperparameter tuning, Algorithm 2 reveals a smoother post-training relationship between generalization and distance to the origin, offering a more efficient alternative.
@@
We also compare LLPF with AutoNEB pmlr-v80-draxler18a , which bends linear interpolation between low-loss models to form a low-loss path. Our proposed LLPF algorithms differ from AutoNEB in several key aspects. First, unlike AutoNEB, which often finds paths with low testing loss, LLPF does not guarantee this, highlighting an inherent difference in the resulting low-loss paths. Second, LLPF-generated paths exhibit consistent properties, such as test loss and L2 distance to the origin, across different simulations, as shown by the narrow shaded regions in Figure 3 and Figure 3, whereas AutoNEB paths are often inconsistent, as shown in Figure 7 of  pmlr-v80-draxler18a . Third, AutoNEB does not explore regions near the origin, while Algorithm 2 explicitly targets this space. Fourth, our algorithm can be configured to explore low-loss spaces with even lower loss than those found by AutoNEB. This result is presented in Appendix A.7.
The result in Appendix A.7 also provides quantitative support for the continuity of the low-loss space. Since models can have much smaller loss values than low-loss paths, our ability to construct a path with a maximum loss lower than AutoNEB’s minimum loss offers stronger empirical evidence that continuous low-loss paths exist between low-loss models.